Calculate T ; N; B; "\kappa" ; "\tau" of the curve x(t) = (t; t 2 ; t 4) at the point (1; 1; 1).


1
Expert's answer
2022-06-26T17:19:42-0400
"\vec{x}(t)=\langle t,t^2,t^4\rangle"

"\vec{x}'(t)=\langle 1,2t,4t^3\rangle"

"|\vec{x}'(t)|=\sqrt{(1)^2+(2t)^2+(4t^3)^2}"

"\vec{x}'(1)=\langle 1,2,4\rangle"

"|\vec{x}'(1)|=\sqrt{1^2+2^2+4^2}=\sqrt{21}"

a.


"\vec{T}(t)=\dfrac{\vec{x}'(t)}{|\vec{x}'(t)|}"

"\vec{T}(1)=\dfrac{\vec{x}'(1)}{|\vec{x}'(1)|}=\langle \dfrac{1}{\sqrt{21}},\dfrac{2}{\sqrt{21}},\dfrac{4}{\sqrt{21}}\rangle"



b.


"\vec{N}(t)=\dfrac{\vec{T}'(t)}{|\vec{T}'(t)|}"

"\vec{T}'(t)=\langle 0,2,12t^2\rangle"

"|\vec{T}'(t)|=\sqrt{(0)^2+(2)^2+(12t^2)^2}"

"\vec{T}'(1)=\langle 0,2,12\rangle"

"|\vec{T}'(1)|=\sqrt{(0)^2+(2)^2+(12)^2}=2\sqrt{37}"

"\vec{N}(1)=\dfrac{\vec{T}'(1)}{|\vec{T}'(1)|}=\langle 0,\dfrac{1}{\sqrt{37}},\dfrac{6}{\sqrt{37}}\rangle"

c.


"\vec{B}(t)=\vec{T}(t)\times \vec{N}(t)"

"\vec{B}(1)=\vec{T}(1)\times \vec{N}(1)=\begin{vmatrix}
 \vec{i} & \vec{j} & \vec{k} \\
 \dfrac{1}{\sqrt{21}} & \dfrac{2}{\sqrt{21}} & \dfrac{4}{\sqrt{21}} \\
 0 & \dfrac{1}{\sqrt{37}} & \dfrac{6}{\sqrt{37}} 
\end{vmatrix}"

"= \dfrac{8}{\sqrt{777}}\vec{i}- \dfrac{6}{\sqrt{777}}\vec{j} + \dfrac{1}{\sqrt{777}}\vec{k}"

"\vec{B}(1)=\langle \dfrac{8}{\sqrt{777}},- \dfrac{6}{\sqrt{777}}, \dfrac{1}{\sqrt{777}}\rangle"

d.


"\kappa=\dfrac{|\vec{T}'(t)|}{|\vec{x}'(t)|}"

"t=1"

"\kappa=\dfrac{|\vec{T}'(1)|}{|\vec{x}'(1)|}=\dfrac{2\sqrt{37}}{\sqrt{21}}"


e.


"\vec{x}''(t)=\langle 0,2,12t^2\rangle"

"\vec{x}''(1)=\langle 0,2,12\rangle"

"\vec{x'}(1)\times \vec{x''}(1)=\begin{vmatrix}
 \vec{i} & \vec{j} & \vec{k} \\
 1 & 2& 4\\
 0 & 2 & 12 
\end{vmatrix}"

"=16\vec{i}- 12\vec{j} +2\vec{k}"

"\vec{x'}(1)\times \vec{x''}(1)=\langle 16,-12,2\rangle"

"|\vec{x'}(1)\times \vec{x''}(1)|=\sqrt{(16)^+(-12)^2+(2)^2}=2\sqrt{101}"



"\vec{x}'''(t)=\langle 0,0,24t\rangle"

"\vec{x}'''(1)=\langle 0,0,24\rangle"

"(\vec{x'}(1)\times \vec{x''}(1))\cdot\vec{x}'''(1)=0-0+48=48"

"\tau=\dfrac{(\vec{x'}(t)\times \vec{x''}(t))\cdot\vec{x}'''(t)}{|\vec{x'}(t)\times \vec{x''}(t)|^2}"

"t=1"


"\tau=\dfrac{(\vec{x'}(1)\times \vec{x''}(1))\cdot\vec{x}'''(1)}{|\vec{x'}(1)\times \vec{x''}(1)|^2}"

"=\dfrac{48}{404}=\dfrac{12}{101}"